# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

### v2

@@ Abstract
Spiking Neural Networks (SNNs), as a biologically plausible alternative to Artificial Neural Networks (ANNs), have demonstrated advantages in terms of energy efficiency, temporal processing, and biological plausibility. However, SNNs are highly sensitive to distribution shifts, which can significantly degrade their performance in real-world scenarios. Traditional test-time adaptation (TTA) methods designed for ANNs often fail to address the unique computational dynamics of SNNs, such as sparsity and temporal spiking behavior. To address these challenges, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method specifically designed for SNNs. SPACE leverages the inherent spike dynamics of SNNs to maximize the consistency of spike-behavior-based local feature maps across augmented versions of a single test sample, enabling robust adaptation without requiring source data. We evaluate SPACE on multiple datasets. Furthermore, SPACE exhibits robust generalization across diverse network architectures, consistently enhancing the performance of SNNs on CNNs (such as VGG and ResNet), Transformer models, and ConvLSTM architectures. Experimental results show that SPACE outperforms state-of-the-art methods, highlighting its effectiveness and robustness in real-world settings.
@@ 1 Introduction
Recent advancements in neuroscience-inspired computing have placed Spiking Neural Networks (SNNs) at the forefront of research as a biologically plausible alternative to traditional Artificial Neural Networks (ANNs). While ANNs have achieved remarkable success in various domains, their reliance on dense, black-box architectures often limits their interpretability and energy efficiency. In contrast, SNNs emulate the sparse, event-driven dynamics of biological neurons, providing several advantages in terms of computational efficiency, temporal processing, and explainability [4, 12, 37]. However, the unique features that make SNNs advantageous—such as their reliance on temporal information and sparse spiking—also pose significant challenges in dynamic real-world environments. SNNs encode information dynamically across time, making them highly sensitive to changes in input data distribution [2]. This sensitivity can result in drastic changes to spiking behavior across the network, ultimately degrading performance when faced with distribution shifts. Understanding and enhancing the robustness of SNNs under such conditions is therefore a critical research direction.
In real-world applications, it is common that the distribution of the training data (a.k.a. source domain) and the test data (a.k.a. target domain) are different due to variations in lighting, background, sensor noise, or environmental conditions. This phenomenon, known as domain shift, can significantly degrade the performance of machine learning models [15, 25]. SNNs are particularly vulnerable to domain shifts, as even minor changes in input data can disrupt the temporal dynamics and spiking behavior in intermediate layers, leading to poor generalization. Traditional test-time adaptation (TTA) solutions such as retraining or fine-tuning on target domain data with help of source data related proxy task [41], statistics of source data [33], or a batch of test data [42], are often impractical in real-world scenarios due to the unavailability of source data or other test samples, particularly in privacy-sensitive or resource-constrained environments. Source-free and single-instance TTA methods [9, 23, 47] offer a promising alternative by enabling the model to adapt to the target domain during inference without access to source data, which is typically defined (and also in this work) as one iteration of training adaptation with one test sample.
Among existing source-free and single-instance TTA methods, MEMO [47] and SITA [23] are two representative approaches, but both face fundamental limitations when applied to SNNs. MEMO optimizes the entropy of the averaged prediction across multiple augmented views of a test sample, enforcing output probability consistency. However, this strategy may not be optimal for SNNs. In SNNs, output probability distributions reflect only high-level semantic information and fail to capture the fine-grained spiking dynamics that encode essential features which prevents MEMO from effectively mitigating domain shifts within SNNs’ internal representations. On the other hand, SITA adapts batch normalization (BN) statistics by shifting them toward those computed from test-time augmentations, but this approach is also unsuitable for SNNs. In SNNs, many models lack BN layers [27, 46], which limits the applicability of BN-based adaptation methods such as SITA in SNNs. Additionally, due to the sparsity of SNN activations, modifying BN statistics does not significantly alter neuronal spiking patterns, further reducing its ability to facilitate effective adaptation.
@@
In this work, we propose a novel method names SPike-Aware Consistency Enhancement (SPACE), which is the first source-free and single-instance TTA approach specifically designed for SNNs. SPACE performs adaptation using only a single test point without access to source data, making it particularly suitable for real-world scenarios where source data are unavailable or privacy-sensitive. By leveraging the inherent spike dynamics of SNNs, SPACE maximizes the consistency in spike-behavior-based local feature maps across augmented samples of the same input, ensuring robust and efficient adaptation in deep SNNs. Unlike existing TTA methods, which often overlook the unique characteristics of SNNs, SPACE directly exploits spike-based representations for adaptation. Our contributions can be summarized as follows
@@
    To the best of our knowledge, we are the first TTA method tailored for SNNs, which operates effectively using only a single test sample. This addresses the unique challenges of adapting SNNs in scenarios where source data are unavailable, while maintaining efficiency and scalability.
@@
    Our method introduces a consistency-driven optimization framework to enhance the similarity of spike-behavior-based local feature maps across augmented samples. By leveraging spike dynamics, this approach ensures robust adaptation in deep SNNs.
@@
    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [26], Tiny-ImageNet, ImageNet [6], and the neuromorphic dataset DVS Gesture [1]. Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on SNN-VGG [24, 40], SNN-ResNet11 [14, 27], Spike-driven Transformer V3 [44], and SNN-ConvLSTM [39]. The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
@@ 2.1 SNN-based Deep Learning
Over the years, significant progress has been made in SNN-based deep learning. Many approaches primarily relied on converting pre-trained ANN models into SNNs, enabling SNNs to inherit the representational power of ANNs while benefiting from spike-based computations [3, 8, 19, 22]. However, such conversion methods often suffer from performance degradation due to differences in activation dynamics. To overcome this, direct training of SNNs using surrogate gradient methods has gained traction, enabling end-to-end optimization of spiking models [7, 11, 30, 32, 46]. Rathi et al. [34] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance. These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [27, 43] and spiking recurrent networks [45, 48], which have denoted promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
A critical limitation of current SNN-based deep learning approaches is their poor robustness to testing-time variations. Unlike ANNs, which operate on continuous activations and can leverage well-established testing-time adaptation techniques, SNNs are highly sensitive to input perturbations due to their dependence on precise spike timing and temporal dynamics. According to our experimental results presented in section 4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time. Addressing these shortcomings is essential for deploying SNNs in practical applications, particularly those requiring real-time adaptability.
@@ 2.2 Test-time Adaptation
Although these approaches eliminate the need for source data, they often assume batch-level adaptation, relying on test-time statistics computed over multiple samples. This assumption limits their applicability in real-world settings where only single test points are available at a time. MEMO [47] addresses this scenario, proposing minimizing the marginal output distribution in the increases of the single test point. The objective in MEMO encourages the model to produce consistent predictions across different augmentations, thus enforcing the invariances encoded in these augmentations, while also maintaining confidence in its predictions. Another TTA method, SITA [23], adapts batch normalization statistics at inference by shifting them toward those computed from a batch of augmented versions of the single test instance.
MEMO has shown that enforcing consistency among augmented samples improves TTA performance. However, in deep SNNs, where information is transmitted through discrete spikes, backpropagation tied to output objectives suffers from temporal information loss, making effective TTA challenging. Moreover, MEMO fails to capture the domain shift-induced variations in the internal behavior of the SNN, such as changes in neuron activations and spike patterns, which are critical for robust adaptation in SNNs. Although SITA avoids backpropagation, it also struggles to accurately model the intrinsic temporal dynamics of SNNs, limiting its ability to align augmented samples effectively. To overcome these issues, we enforce consistency by directly aligning the spike patterns and behaviors of augmented samples within the SNN framework. Unlike BN-dependent methods like SITA, our approach works regardless of the presence of BN layers, ensuring broader compatibility with various SNN architectures.
@@ 3.1 Preliminary
SNNs emulate biological neurons by transmitting and encoding information through discrete spikes. Among various neuron models, the Leaky Integrate-and-Fire (LIF) neuron [5] is widely adopted for its balance between biological realism and computational efficiency. Its membrane potential dynamics follow
@@
Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for weight connections between pre- and post-neurons. Figure 1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
@@
While this is not an exact solution, it is valid because a reset occurs after a spike is generated when $U\geq U_{th}$.
These foundational techniques for SNN training, including LIF dynamics and surrogate gradient approximations, form the basis for our proposed method, which focuses on enhancing TTA by ensuring consistency in spiking behavior across augmented samples.
@@ 3.2 SPike-Aware Consistency Enhancement
The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Figure 2. We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample. This augmented batch is passed through the model to obtain local feature maps, represented by spike counts over time. Next, the model is adapted by maximizing the similarity across the feature maps of the augmented samples, promoting consistency in feature representations. Finally, the adapted model predicts the label of the original test sample, ensuring robust performance under test-time conditions. This approach enhances the model’s ability to generalize to unseen test samples, particularly in shifted domains. The overall method is presented in Algorithm 1, and the details are introduced below. The code will be made available after acceptance.
@@
SPACE aims to adapt pre-trained SNN-based models $M_{\theta}$ (with parameters $\theta\in\Theta$) during inference to improve performance under distribution shifts, without requiring ground truth labels or access to source data. No specific training process is required, nor do we impose particular constraints on the model. The only assumptions are that $\theta$ can be adjusted and that the model produces intermediate feature maps $\mathbf{F}(\mathbf{x})$, which is differentiable with respect to $\theta$ and can be utilized for further analysis and adaptation. Here, $\mathbf{x}\in\mathcal{X}$ is a single test point that is presented to $M_{\theta}$. It is worth noting that this is a reasonable assumption, supported by a recent advance [36] that suggest gradient back-propagation can be implemented on spiking neuromorphic hardware. Aligned with previous work [23, 47], we achieve test-time robustness by leveraging data augmentations and a self-supervised adaptation objective. In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples =$\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$. Using this batch, we align the spike dynamics extracted from augmentations, ensuring that the model produces reliable and robust predictions in the presence of distribution shifts.
Different from MEMO [47], which adapts the model during test time via the conditional output distribution $p_{\theta}(y|\mathbf{x})$ of the model’s last layer, SPACE leverages the dynamic behavior of intermediate layers in the SNN that plays a crucial role in determining the temporal information and semantic representation of the final output. SNNs encode information through temporal dynamics and spike sparsity. These behavioral characteristics are particularly prominent in the intermediate layers and can be represented by feature maps $\mathbf{F}(\mathbf{x})$. Unlike ANNs, these characteristics in SNNs cannot be fully captured by the output probability distribution alone. Another advantage of intermediate feature maps is its highly sensitivity to changes in the input distribution due to the event-driven nature of SNNs. When the input data undergoes a distribution shift, the spiking patterns in the intermediate layers may change significantly, thereby impacting the final output. Optimizing the consistency of spiking behavior in intermediate layers can effectively mitigate feature variations caused by distribution shifts, thereby enhancing the robustness of models.
@@
To leverage the feature maps from the intermediate layers of the SNN, we focus on the spike activity over the entire time window. As described in subsection 3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$. The temporal spike patterns of the network are captured in the feature map $\mathbf{O}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the number of time steps, $C$ represents the number of channels, and $H$ and $W$ are the spatial dimensions of the local feature maps. $\mathbf{x}_{i}$ refers to the $i$-th sample in the augmented batch. Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network, depending on whether the final feature map is pooled into a 1$\times$1 matrix. Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{2}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
There are several reasons for choosing the total spike counts as the observation metric. First, SNNs typically operate over a time window ranging from 25 to thousands time steps [4, 13, 27, 34, 38], meaning that spike patterns across many time steps may be identical due to the repetitive nature of the spike dynamics. By comparing the cumulative spike counts instead of individual time-step patterns, we can effectively capture the temporal accumulation of spike activity while reducing redundancy. Second, the total spike counts reflects differences between augmented samples based on their overall spike dynamics, which is essential for our goal of aligning feature maps across augmentations. Finally, this representation significantly reduces the computational cost, as it condenses the temporal information into an aggregate value without requiring a step-by-step comparison of spike patterns.
@@
1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, model $\theta$, learning rate $\eta$
2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the updated model
@@
9:  Pass $\mathcal{B}$ through SNN, obtain =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
@@
15:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Equation 5, update $\theta$ via SGD with $\eta$
@@
17:  $y^{*}=\arg\max_{y}p_{\theta^{*}}(y|\mathbf{x})$
18:  Return: $y^{*}$

### v3

@@ Abstract
Spiking Neural Networks (SNNs), as a biologically plausible alternative to Artificial Neural Networks (ANNs), have demonstrated advantages in terms of energy efficiency, temporal processing, and biological plausibility. However, SNNs are highly sensitive to distribution shifts, which can significantly degrade their performance in real-world scenarios. Traditional test-time adaptation (TTA) methods designed for ANNs often fail to address the unique computational dynamics of SNNs, such as sparsity and temporal spiking behavior. To address these challenges, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method specifically designed for SNNs. SPACE leverages the inherent spike dynamics of SNNs to maximize the consistency of spike-behavior-based local feature maps across augmented versions of a single test sample, enabling robust adaptation without requiring source data. We evaluate SPACE on multiple datasets. Furthermore, SPACE exhibits robust generalization across diverse network architectures, consistently enhancing the performance of SNNs on CNNs, Transformer, and ConvLSTM architectures. Experimental results show that SPACE outperforms state-of-the-art ANN methods while maintaining lower computational cost, highlighting its effectiveness and robustness for SNNs in real-world settings. The code will be available at https://github.com/ethanxyluo/SPACE.
@@ 1 Introduction
Recent advancements in neuroscience-inspired computing have placed Spiking Neural Networks (SNNs) at the forefront as a biologically plausible alternative to traditional Artificial Neural Networks (ANNs). While ANNs have achieved remarkable success across domains, their reliance on dense, black-box architectures often limits interpretability and energy efficiency. In contrast, SNNs emulate the sparse, event-driven dynamics of biological neurons, offering advantages in computational efficiency, temporal processing, and explainability [13, 4, 44]. However, the very features that make SNNs appealing—temporal coding and sparse spiking—also introduce fragility in dynamic real-world environments. Because SNNs encode information over time, they can be highly sensitive to shifts in the input distribution [2], which may induce substantial changes in layer-wise spiking activity and degrade downstream performance. At the same time, the historical SNN–ANN accuracy gap on standard vision benchmarks has narrowed considerably [58, 52], foregrounding generalization and robustness as primary objectives for practical SNN deployment [10, 51]. In application domains central to SNNs—event-driven sensing and low-power edge inference—the ability to remain robust to distribution shift is thus a deployment prerequisite rather than a post-hoc enhancement.
In real-world deployments, it is common for the training (a.k.a. source) and test (a.k.a. target) distributions to differ owing to changes in lighting, background, sensor noise, or environmental conditions. Such domain shifts are known to impair model performance [18, 30]. Prior studies have specifically noted robustness challenges of SNNs under distribution shift [38, 27, 23], but remain hard to deploy online at scale. Mei et al. [38] perform gradual, unsupervised adaptation via teacher–student self-training across intermediate domains, which requires iterative weight updates, pseudo-label calibration, and multiple passes during deployment—assumptions that clash with single-sample, low-latency edge constraints. Karilanova et al. [27] address only temporal-resolution shifts via parameter remapping that presumes a known source–target sampling ratio and synchronized binning, offering no per-instance online adaptation and limited coverage of broader appearance/sensor shifts. To address the limitations, we adopt SNN-specific test-time adaptation (TTA) to meet practical deployment constraints.
Existing TTA approaches often rely on source data or batched target samples—e.g., using source-related proxy tasks [48], source statistics [40], or optimization over a batch of test inputs [49, 54, 34]. Such assumptions are frequently impractical due to privacy, regulatory, and on-device resource constraints. We therefore focus on the source-free, single-instance regime, where the model adapts online using only the current test sample and without access to source data or target batches. Note that several influential ANN-oriented TTA methods fall outside this regime. For example, SHOT [37] requires mini-batches of target data and multi-pass self-training in a static feature space while freezing a standalone classifier head; in SNNs, classification is tightly coupled to time-evolving spiking dynamics, leaving no stationary head to freeze and rendering distance-based pseudo-labeling insensitive to temporal structure. Similarly, TAST [25] relies on class-prior regularization and batch-level target statistics and is typically optimized over target mini-batches—assumptions misaligned with sparse, discrete spike trains and incompatible with the single-instance constraint.
Within the truly source-free, single-instance family, MEMO [56] and SITA [28] are representative but face fundamental limitations for SNNs. MEMO enforces output-consistency by minimizing the entropy of the averaged prediction over augmented views; however, output probabilities capture only coarse semantics and provide weak control over the fine-grained temporal spiking dynamics that govern internal representations. SITA adapts batch-normalization (BN) statistics using test-time augmentations, yet many SNN architectures omit BN layers [55, 32]; even when present, sparse activations limit the effect of BN-moment shifts on spike generation. These gaps motivate an SNN-specific TTA framework that explicitly engages temporal spiking dynamics, avoids reliance on BN, and respects the source-free, single-instance constraints to improve robustness under domain shift.
@@
In this work, we propose a novel method named SPike-Aware Consistency Enhancement (SPACE), which is the first source-free and single-instance TTA approach specifically designed for SNNs. SPACE performs adaptation using only a single test point without access to source data, making it particularly suitable for real-world scenarios where source data are unavailable or privacy-sensitive. By leveraging the inherent spike dynamics of SNNs, SPACE maximizes the consistency in spike-behavior-based local feature maps across augmented samples of the same input, ensuring robust and efficient adaptation in deep SNNs. Unlike existing TTA methods, which often overlook the unique characteristics of SNNs and the demanding of single instance scenarios, SPACE directly exploits spike-based representations for adaptation. Our contributions can be summarized as follows
@@
    To the best of our knowledge, we are the first TTA method tailored for SNNs, which operates using only a single test sample. This addresses the unique challenges of adapting SNNs in scenarios where source data are unavailable, while maintaining efficiency and accuracy.
@@
    Our method introduces a consistency-driven optimization framework to enhance the similarity of spike-behavior-based local feature maps across augmented samples of the single point. By leveraging spike dynamics, this approach ensures robust adaptation in deep SNNs.
@@
    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [31], Tiny-ImageNet, ImageNet [7], and the neuromorphic dataset DVS Gesture [1]. Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on SNN-VGG [47, 29], SNN-ResNet [17, 32], Spike-driven Transformer V3 [52], and SNN-ConvLSTM [46]. The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
@@ 2.1 SNN-based Deep Learning
Over the years, significant progress has been made in SNN-based deep learning. Many approaches primarily relied on converting pre-trained ANN models into SNNs, enabling SNNs to inherit the representational power of ANNs while benefiting from spike-based computations [9, 3, 22, 26]. However, such conversion methods often suffer from performance degradation due to differences in activation dynamics. To overcome this, direct training of SNNs using surrogate gradient methods has gained traction, enabling end-to-end optimization of spiking models [55, 39, 12, 8, 36]. Rathi et al. [41] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance. These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [50, 32], spiking recurrent networks [53, 57], and spiking transformer networks [58, 52], which have denoted promising results in diverse tasks like image classification, speech recognition, and event-driven sensor data analysis.
@@ 2.2 Test-time Adaptation
Although these methods avoid source data, many assume batch-level adaptation, relying on test-time statistics over multiple samples—an impractical requirement when only a single test point is available. MEMO [56] addresses the single-instance setting by minimizing entropy and enforcing prediction consistency across augmentations, thereby promoting augmentation-invariant semantics but offering limited control over temporal spiking dynamics under shift. SITA [28] adapts batch-normalization statistics using augmented views of the same instance; however, many SNNs lack BN [55, 32], and sparse activations diminish the effectiveness of moment updates, especially under severe corruptions. To overcome these limitations, we align spike patterns and temporal features across augmentations without relying on BN, yielding an SNN-specific TTA mechanism compatible with diverse architectures and truly single-sample operation, while offering a more stable and informative alignment signal for robustness.
2.3 Generalization of SNNs
Recent state-of-the-art SNNs [58, 52] achieve performance on clean data comparable to strong ANNs, shifting the research focus from in-distribution accuracy to generalization and robustness under real-world shifts. Several works have begun to examine these issues in SNNs [10, 15, 51]. Guo et al. [15] study Source-Free Domain Adaptation (SFDA) in a batch/epoch setting, where the model is updated after processing target-domain batches and typically assumes access to a pre-collected target dataset. In contrast, our work tackles TTA in a stricter, online regime: we adapt to each incoming sample individually and on the fly, without labels, source data, or target-set pre-collection. This per-sample update schedule reduces latency and memory footprint, making the approach directly applicable to streaming, autonomous deployments where SNNs are expected to excel.
@@ 3.1 Preliminary
SSNNs emulate biological neurons by transmitting and encoding information through discrete spikes. Among various neuron models, the Leaky Integrate-and-Fire (LIF) neuron [6] is widely adopted for its balance between biological realism and computational efficiency. Its membrane potential dynamics follow
@@
Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for connections between pre- and post-neurons. Figure 1 shows the LIF model, depicting membrane potential evolution and spike generation in response to input spikes [11].
@@
While this is not an exact analytical solution, it is nevertheless valid because a reset promptly occurs after a spike is generated when $U\geq U_{th}$.
These SNN foundational techniques underpin our method to enhance TTA by ensuring consistent spiking behavior across augmented samples.
@@ 3.2 SPike-Aware Consistency Enhancement
The proposed algorithm through spike-aware consistency enhancement (SPACE) for TTA is designed to improve test-time robustness of SNNs, as shown in Figure 2. We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample. This augmented batch is passed through the model to obtain local feature maps, represented by spike counts over time. Next, the model is adapted by maximizing the similarity across the feature maps of the augmented samples, promoting consistency in feature representations. Finally, the adapted model predicts the label of the original test sample, ensuring robust performance under test-time conditions. This approach enhances the model’s ability to generalize to unseen test samples, particularly in shifted domains. The overall method is presented in Algorithm 1, and the details are introduced below.
@@
SPACE aims to adapt pre-trained SNN-based models $M_{\theta}$ (with parameters $\theta\in\Theta$) during inference to improve performance under distribution shifts, without requiring ground truth labels or access to source data. No specific training process is required, nor do we impose particular constraints on the model. The only assumptions are that $\theta$ can be adjusted and that the model produces intermediate feature maps $\mathbf{F}(\mathbf{x})$, which is differentiable with respect to $\theta$ and can be utilized for further analysis and adaptation. Here, $\mathbf{x}\in\mathcal{X}$ is a single test point that is presented to $M_{\theta}$. It is worth noting that this is a reasonable assumption, supported by a recent advance [43] that suggest gradient back-propagation can be implemented on spiking neuromorphic hardware. Concretely, we decompose the SNN into a spike-based feature extractor and a classifier, writing $M_{\theta}(\mathbf{x})=C_{\theta_{C}}\!\big(E_{\theta_{E}}(\mathbf{x})\big)$ with $\theta=(\theta_{E},\theta_{C})$. Here, the extractor $E_{\theta_{E}}$ denotes the main body of the network (e.g., CNN/Transformer/ConvLSTM blocks), and the classifier $C_{\theta_{C}}$ consists of the last one or two readout layers.
Aligned with previous work [28, 56], we achieve test-time robustness by leveraging data augmentations and a self-supervised adaptation objective. In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples $\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$. Using this batch, we align the spike dynamics extracted from augmentations, ensuring that the model produces reliable and robust predictions in the presence of distribution shifts. SPACE differs from MEMO [56] by operating not on the conditional output $p_{\theta}(y|\mathbf{x})$ but on the spike dynamics of intermediate layers that determine temporal coding and semantic representations. Specifically, we observe the feature map $\mathbf{O}(\cdot)\in\mathbb{R}^{T\times C\times H\times W}$ at the deepest layer of $E_{\theta_{E}}$ that still retains spatial support. Here $T$ denotes the number of time steps, $C$ represents the number of channels, $H$ and $W$ are the spatial dimensions of the local feature maps. If the last pre-classifier feature map has shape with $H\times W>1$, we use this layer; if global pooling or stride reduces it to $1\times 1$, we use the penultimate layer of $E_{\theta_{E}}$. This choice balances semantic abstraction and alignment granularity—using a $1\times 1$ map yields too few alignment targets and risks overfitting, whereas too shallow a layer may overemphasize low-level variations.
@@
To leverage the intermediate spike activity over the entire time window, let $\mathbf{O}(\mathbf{x}_{i})\in\{0,1\}$ denote binary spikes (1=fire, 0=silent, refers to subsection 3.1) at the chosen extractor layer for an augmented view $\mathbf{x}_{i}$. We aggregate spikes over time to form a spike-aware feature maps collection $\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{2}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
We choose total spike counts for three main reasons: 1) SNNs often run for tens to thousands of steps [45, 16, 32, 41], making stepwise matching redundant and costly. 2) Aggregation smooths the loss landscape: if spikes jitter in time but counts match, the loss remains unchanged, so gradients issue a clear directive—make this neuron more or less excitable for this input—producing stable, macro-level alignment of core feature activations across the augmented views. 3) distribution shift primarily manifests as changes in spiking rates and spatial activation support in intermediate layers; $\mathbf{F}(\cdot)$ provides a direct, informative summary of these changes that output probabilities cannot capture.
@@
1:Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, SNN model with extractor parameters $\theta_{E}$, and learning rate $\eta$
2:Output: Prediction $y^{*}$ via updated model
@@
9: Pass $\mathcal{B}$ through the SNN feature extractor, obtain $\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
@@
15: Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Equation 5, update $\theta_{E}$ via SGD with learning rate $\eta$
@@
[…]
18:Return: Prediction result $y^{*}$